OpenStudy (anonymous):

how would use the substitution method to solve the following system of equations. Provide the solution to the system. 2x + y = 5 3x – 2y = 25

OpenStudy (agent47):

solve for y in the first equation: y=5-2x plug the value of y into the 2nd equation: 3x-2(5-2x)=25

OpenStudy (agent47):

Solve for x in the 2nd equation: 3x-10+4x=25 7x=35 x=5

OpenStudy (agent47):

Plug the value of x back into the first equation and solve for y: y=5-2x y=5-2*5 y=5-10=-5